You are here: Foswiki>Commissioning Web>Com20080611 (11 Jun 2008, AndrewRakich)Edit Attach

Commissioning Notes from 10th-11th June 2008 MST

Observers: none (at LBT) D. Miller, J. Hill, A. Rakich (in Tucson).
OSA: D. Gonzalez Huerta

Summary

Just started getting stuff working when 60 mi/hr winds kicked up and shut us down for the rest of the night. Were able to confirm that the guiding offset problem had been corrected on-axis, and to identify a new problem noted in issuetrak 1432 to do with a PCS problem with guiding.

Night time Activities

(Times in MST below)

07:50 Open Dome, reset encoders. T=13degC D=-7degC Wind: 5-15 m/sec from SW.

7:50-8:30 SX M1 cell confused about voltage state, fix: turn it off and turn it on again.

9:24 send preset, rotator trouble.

9:30 somthing not working in new implementation of GCS, library path has changed with new build and now IDL routines have problems. (later, fixed now).

9:35 tracking star but nothing on either IRTC or guider, going to brighter star and spiral searching. Going to WT10-367

9:44 see a diffraction spike on guide camera, went to IRTC and spiral search, find the star at about 30" off-axis.

9:54 going back to Pearson star for guiding and active optics BS9162

10:05 collimating then had to stop for 60 mi/hr wind, stayed shut all night.

3:35 submitted Issuetrak 1432 to do with PCS not correcting consistant x-error in guiding.

3:51 Doug suddenly lost all connections to mountain apart from AGW server. David reports PMC is not running at all. Is there still some cronjob chewing up time?

-- DavidThompson - 10 Jun 2008
Topic revision: r4 - 11 Jun 2008, AndrewRakich
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback